Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability Question

Expert replies
by sukhman » Sun Sep 08, 2013 11:53 am
If five fair coins are tossed together , what is the probability of getting at most 3 tails .
Answer 13/16
Join the discussion
Source: — Problem Solving |

by Java_85 » Sun Sep 08, 2013 12:42 pm
We have 5 coins and each can be tail or head == > 5^2 = 32 possible states
getting at most 3 tails means 1, 2 or 3 tails, in other words 4 or 5 heads.
how many states with 5 heads ==>1
how many states with 4 heads ==>5 |=> 5+1 =6 ==> 32-6=26 states with 1, 2 or 3 tails,

Probability = 26/32= 13/16
Join the discussion

by sanjoy18 » Sun Sep 08, 2013 12:44 pm
probability of getting at most 3 tails= 1- probability of getting at least 4 tails

probability of getting at least 4 tails
=probability of getting 4 tails + probability of getting 5 tails
= 5/2^5 +1/2^5
= 6/32

hence required probability= 1-6/32= 26/32= 13/16
[NOTE: probability of getting 4 tails..their would be 5 cases TTTTH,TTTHT,TTHTT,THTTT,HTTTT]
Join the discussion

by sanjoy18 » Sun Sep 08, 2013 12:52 pm
This problem can be solved by using binomial distribution formula

p(x)=(ncx)*p^x q^(1-x)
where p =probability getting tail=1/2
q=1-p=1/2

according to the problem, we have to find out
P(x<=3)=p(x=0)+p(x=1)+p(x=2)+p(x=3)-----------------(A)
=(1/32)*[5c0+5c1+5c2+5c3]
=26/32=13/16 (ANS)

(A) can be written as
p(x<=3)=1-[p(x=4)+p(x=5)]
=1-6/32
=13/16
Join the discussion